How much water should be added to 1 gallon of antifreeze to obtain a solution that is 60% antifreeze?
Answer by Jk22(389)

1 gallon should equal 60%. Let w water gallon added : the total is 1+w gallon, of which 60% is antefreeze, hence 1 gallon :

0.6*(1+w)=1=0.6+0.6w

0.4=0.6w

w=4/6=2/3~0.66 gallons.

Verif : Total : 1+2/3, % antifreeze : 1/(1+2/3)=3/(3+2)=3/5=60/100.